OpenStudy (scienceteen8):

Bill has 2/3 of his pizza left from lunch. If he eats 3/7 more of the pizza, how much pizza is left? A.1/4 B.1/3 C.1/21 D.5/21

OpenStudy (campbell_st):

are you sure you have the answer choices correct

OpenStudy (scienceteen8):

yes

OpenStudy (campbell_st):

ok... so you need a common denominator... what would i be..?

OpenStudy (campbell_st):

2/3 and 3/7 what is the common denominator..?

OpenStudy (scienceteen8):

i multiply 3 and 7 right

OpenStudy (campbell_st):

correct.... so the common denominator is 21 so 2/3 = ?/21 what's the value of ?

OpenStudy (scienceteen8):

6/21?

OpenStudy (campbell_st):

yo also need to change the remaining pizza 4/7 to ?/21

OpenStudy (campbell_st):

well look at its this way \[\frac{2}{3} \times \frac{7}{7} = \frac{?}{21}\]

OpenStudy (scienceteen8):

im going to go with 5/21

OpenStudy (campbell_st):

well no..... if you eat 3/7 how many 7ths are remaining...?

OpenStudy (scienceteen8):

4

OpenStudy (campbell_st):

so write 4/7 as a fraction with a denominator of 21

OpenStudy (scienceteen8):

so if there is going to be a denominator with 21 would it be 1/21

OpenStudy (campbell_st):

well its a poorly worded question.... in that it doesn't explicitly say he eats 3/7 of the 2/3 but your previous answer seems to be correct 14/21 - 3/7 or 14/21 - 9/21 = 5/21

OpenStudy (scienceteen8):

ok thank you for the help idk why my school always does this lol